UW CHEMISTRY NMR

Announcments & Equipment Logs

UPDATE ! AV301 – SAMPLE HANDLING AND LOCK WORKING OK

The issue of BSMS not responding, leading to sample handling and lock was a temporary issue, probably caused by a voltage surge.  The issue is resolved and the system is functioning normally now. NMR Facility Manager 2/12/2019

Posted in Hardware, Spectrometer | No Comments

UPDATE ! AV500 – THE TBI PROBE IS BACK

The mechanical issue compromising the 1H match has been fixed in-house and the probehead is back in the magnet. We have replaced the broken 1H match knob with a homemade knob.  I will show a picture below that clearly identifies the new homemade knob. Only for 1H Match, you need to use the following tune/match tool […]

Posted in Hardware, Spectrometer | No Comments

ALERT ! AV300 – SAMPLE LANDING ISSUE – UPDATE

Sample spinning with this instrument is not working and this is also related to the  BSMS system’s observation that sample is ‘missing‘.   We are looking into most plausible causes before we decide to pull the probe out, to minimize the down time with the instrument. Please note that the only alternate probehead that can […]

Posted in Hardware, Spectrometer | No Comments

ALERT ! AV500 – PROBE CHANGED TO HCP-Z gradient

We discovered a broken part in the TBI probe that made the 1H matching process untenable. While we are working on this, we have placed the triple resonance 1H{13C, 31P}-Zgradient probe in the magnet, as a temporary measure (hopefully).  We have every intention of returning to the excellent TBI probe once it is ready to […]

Posted in Hardware, Spectrometer | No Comments

ALERT ! AV500 – 1H MATCHING ROD MAY NEED FIXING

We are aware of reports from some users that the 1H matching rod is still not moving smoothly.  We will get to this as soon as possible and post an update of the results. NMR Facility Manager 1/29/2019

Posted in Hardware, Spectrometer | No Comments

UPDATE ! AV500 – SYSTEM READY FOR USE

The probe mechanics are fixed to normalcy.  Please go ahead and use the instrument. Do not hesitate to post in the log book if you encounter difficulty. The tuning/matching rods, are near the magnet legs. They are no longer hanging from underneath the probe.  This will prevent accidental loosening of the screws in the probehead […]

Posted in Hardware, Spectrometer | No Comments

ALERT ! AV500 – DOWN FOR MAINTENANCE

The probehead issue I had mentioned last Friday might take more than a quick fix, upon inspection outside the magnet.   I have blocked the instrument until 11 am hoping that this will be done by then. Barring one,  I have retained the existing reservations on the calendar. If the situation changes, I will let you […]

Posted in Hardware, Spectrometer | No Comments

ALERT ! AV500 UPDATE ON PROBEHEAD – PLEASE READ

Tuning and Matching glitch on 1H channel : The matching knob on 1H channel has lost its traction since it is mostly disengaged from the rod that rotates the matching Capacitor inside the probe.  We need to apply a minimally invasive remedy to fix this and we will do this on Monday, the 28th.  But […]

Posted in Hardware, Spectrometer | No Comments

UPDATE ! AV300 – SYSTEM IS AVAILABLE FOR USE AGAIN – PLEASE READ

The network issue has been patched so that the system is functioning normally now.  You shouldn’t see anything different in operating the instrument, but do read the caveat below : Warning : If you attempt to reboot the computer on your own, you will most certainly break the network configuration that is functional now, and […]

Posted in Host Computer | No Comments

ALERT ! AV300 DOWN WITH NETWORK ISSUE

The network communication between the host computer and the NMR console is broken and the instrument is unusable until further notice. We will post updates with progress made as it becomes available. As of now the reservation calendar will be frozen. Existing reservations will be honored when the instrument is back online.  If your slot […]

Posted in Host Computer, Operating System, Software, Spectrometer | No Comments